# BALF: Budgeted Activation-Aware Low-Rank Factorization for Fine-Tuning-Free Model Compression

### v1

@@ 1 Introduction
Here, we propose a unified framework that builds on Wang et al. (2025c) and is compatible with a broad class of layers. Moreover, we introduce a scalable rank allocator that determines how much to compress each layer based on a Lagrangian relaxation. Together, these form BALF, a complete and general fine-tuning-free compression pipeline.
The main contributions of this work can be highlighted in the following points:
@@
  • Extensive evaluation. Across vision models from ResNet-20 (CIFAR-10) to ResNeXt-101 and ViT-B/16 (ImageNet), our method achieves strong accuracy-compression trade-offs in the fine-tuning-free regime.
@@ 2 Related Work
Deep learning model compression has been extensively studied. The principal categories of compression techniques include quantization, which employs low-bit storage and operations (Jacob et al., 2018; Frantar et al., 2023; Gholami et al., 2021); pruning (Hoefler et al., 2021; Frantar & Alistarh, 2022), which involves the elimination of parameters or parameter groups from the model; and factorization, which divides model layers into smaller sub-layers to lower overall memory usage and computational demands (Jaderberg et al., 2014; Ben Noach & Goldberg, 2020). This paper focuses on the factorization approach.
@@
Existing factorization approaches often require fine-tuning and/or compute-intensive search (Idelbayev & Carreira-Perpiñán, 2020; Yu & Bouganis, 2022). The impossibility of running these on common hardware motivates the need for lighter methods. For example, pruning methods that explicitly aim to be efficient in this regard include Narshana et al. (2023); Murti et al. (2023); Chen et al. (2024); Zhang et al. (2024); Wang et al. (2025a), whereas fine-tuning-free factorization methods are less common—for instance, Yu & Bouganis (2022). Liebenwein et al. (2021); Idelbayev & Carreira-Perpiñán (2020) also report some results in the fine-tuning-free regime.
@@ 3 Preliminaries and Notation
Given a matrix ${\bm{A}}\in\mathbb{R}^{M\times N}$, we denote its SVD by ${\bm{A}}={\bm{U}}{\bm{\Sigma}}{\bm{V}}^{T}$, where ${\bm{U}}\in\mathbb{R}^{M\times M}$ and ${\bm{V}}\in\mathbb{R}^{N\times N}$ are orthogonal matrices, and ${\bm{\Sigma}}=\operatorname{diag}(\bm{\sigma};(M,N))$ is a diagonal matrix with its singular values. Another concept of interest is the $P$-truncated SVD, which is defined by taking only the $P$ largest singular values and their associated singular vectors. We define $\mathcal{T}^{\mathrm{SVD}}_{P}\bigl({\bm{A}}\bigr)={\bm{U}}_{:,:P}{\bm{\Sigma}}_{:P,:P}({\bm{V}}_{:,:P})^{T}$, where $\mathcal{T}^{\mathrm{SVD}}_{P}\bigl(\cdot\bigr)$ can be thought of as an operator that projects a matrix to its rank-$P$ approximation via SVD.
@@
Data whitening (Kessy et al., 2018) (also termed data sphering) is a widely used technique in machine learning and statistics. Here, we are interested in a special case called uncentered whitening. This process transforms data so that its second-moment matrix becomes the identity matrix. Formally, we define an uncentered whitening matrix (UWM) ${\bm{M}}\in\mathbb{R}^{I\times I}$ for the data ${\bm{X}}\in\mathbb{R}^{N\times I}$ as a matrix such that ${\bm{M}}^{T}{\bm{X}}^{T}{\bm{X}}{\bm{M}}=N\overline{{\bm{I}}}_{R}$, where $\overline{{\bm{I}}}_{R}\in\mathbb{R}^{I\times I}$ is the identity on the first $R$ coordinates and zero elsewhere. In this case, $R$ is the rank of $\bm{X}$}, which is equal to the dimension of the subspace spanned by the data if we have enough samples. We further restrict the definition by requiring that $\operatorname{row}({\bm{X}})=\operatorname{range}({\bm{M}})$, which in turn forces the last IRI-R columns of $\bm{M}$} to be zero and makes ${\bm{X}}={\bm{X}}{\bm{M}}{\bm{M}}^{+}$; this assumption and these consequences will be used in later results.
@@
We refer to general tensors as $\bm{\mathsfit{X}}$}. When some tensor is known to be a matrix or vector, we may use $\bm{X}$} and $\bm{x}$}, respectively. For matrices, we denote the Moore-Penrose inverse (also termed the pseudoinverse) of $\bm{M}$} by $\bm{M}$$^{+}$, and the square root of the pseudoinverse by ${\bm{M}}^{+1/2}$.
Deep learning frameworks, such as PyTorch (Paszke et al., 2019), usually include “reshape” and “permute” operators, denoted in this work by $\mathsf{reshape}\left({\bm{\mathsfit{X}}};(s_{1},\dots,s_{n})\right)$ and $\mathsf{permute}\left({\bm{\mathsfit{X}}};(\pi_{1},\dots,\pi_{n})\right)$, respectively.
@@
We denote the identity function by $\mathrm{id}$. Moreover, when talking about convolutional layers, $B$ will denote the batch size; $H_{x},W_{x},C_{x}$ the height, width, and channels of the input (when $x=i$) and output (when $x=o$); and $H_{k},W_{k}$ the spatial size of the kernel. In this context, we also refer to $G$ as the number of groups in the convolutional layer (Xie et al., 2016).
@@ 4 Our Framework
$G$ stands for the number of groups. In the special case where $G=1$, we may simply refer to matrices instead of batches containing a single matrix. Note that, in the preceding definition, matrix multiplication is batched, i.e., $G$ independent matrix multiplications are carried out. The necessity of batch semantics in the group dimension will become apparent later. This condition is satisfied by commonly used densely parameterized layers. In what follows, layers are assumed to operate in batched mode, with $B$ denoting the batch size. We refer to $N$ as the outer dimension.
@@
In some architectures (e.g., transformers), linear layers receive inputs with multiple leading independent dimensions—for example, a tensor of shape $(B,L,D)$. In those cases, we flatten the leading dimensions, reshaping the input to $(BL,D)$, and restore the original shape afterward. We now turn our attention to convolutional layers. Although an ungrouped convolution is a special case of a (possibly grouped) convolution, we present them separately for clarity.
@@
where $\mathsf{im2col}:\mathbb{R}^{B\times C_{i}\times H_{i}\times W_{i}}\to\mathbb{R}^{B\times H_{o}W_{o}\times C_{i}H_{k}W_{k}}$ maps a batch of 2D signals with $C_{i}$ channels into a matrix (see Chellapilla et al. (2006)). The PyTorch (Paszke et al., 2019) equivalent is called $\mathsf{unfold}$ [^1] https://docs.pytorch.org/docs/stable/generated/torch.nn.Unfold.html. It encapsulates the convolution hyperparameters—kernel size, stride, padding, and dilation—but not the number of groups. For brevity, we omit the encapsulated hyperparameters from the notation.
On the other hand, a grouped convolution can be viewed as $G$ separate, parallel convolutions. Each one takes a disjoint slice of $C_{i}/G$ input channels and produces $C_{o}/G$ output channels. Concatenating the $G$ outputs recovers the full $C_{o}$ channels (Krizhevsky et al., 2012; Xie et al., 2016). This also fits our framework.
@@
Additionally, Example 2 and Example 3 have trivial extensions to higher-dimensional convolutions, achieved by replacing $\mathsf{im2col}$ with their higher-dimensional counterparts.
@@
In order to simplify the notation, we define $\mathcal{T}^{\mathrm{SVD}}_{P}\bigl({\bm{\mathsfit{W}}}\bigr)\equiv\overline{P}^{-1}\left(\mathcal{T}^{\mathrm{SVD}}_{P}\bigl(\overline{P}({\bm{\mathsfit{W}}})\bigr)\right)$. In other words, we extend the matrix low-rank projection notion to arbitrary tensors that represent weights of $(\overline{O},\overline{I},\overline{P})$-expressible layers. In the case where $G>1$, we decompose each group $\overline{P}({\bm{\mathsfit{W}}})_{g}$ separately, with the constraint that the per-group rank is the same across groups. This latter constraint has practical reasons, which are discussed in Section B.3.1.
@@
$\mathcal{T}^{D}_{P}\bigl(\overline{P}({\bm{\mathsfit{W}}})_{g}\bigr)=\operatorname{arg}\,\underset{\operatorname{rank}(\overline{P}(\widehat{{\bm{\mathsfit{W}}}})_{g})\leq P}{\min}\;D(\overline{P}({\bm{\mathsfit{W}}})_{g},\overline{P}(\widehat{{\bm{\mathsfit{W}}}})_{g})$
where $D$ is some penalty function and $\widehat{{\bm{\mathsfit{W}}}}$ is the approximate tensor. Traditional SVD-based decomposition corresponds to the case where $D$ is the Frobenius norm of the (reshaped) parameters, i.e., it penalizes distance in parameter space (see Equation 1). We use $\mathcal{T}^{D}_{P}\bigl({\bm{\mathsfit{W}}}\bigr)$ to also denote batched computation, where each group’s matrix is decomposed individually, and omit $\overline{P}$ for ease of notation (as before).
@@ 4.1 Activation-Aware Low-Rank Projections
In order to answer that question, we first need to define what “best” should mean. Without loss of generality, fix a group $g$. Clearly, $\operatorname{rank}\left(\overline{I}({\bm{\mathsfit{X}}})\mathcal{T}_{P}\bigl({\bm{\mathsfit{W}}}\bigr)\right)\leq P$. Moreover, by the EYM theorem, the best rank-$P$ approximation of the output of $f$ in terms of the Frobenius norm is $\mathcal{T}^{\mathrm{SVD}}_{P}\bigl(\overline{I}({\bm{\mathsfit{X}}})\overline{P}({\bm{\mathsfit{W}}})\bigr)$. Hence, a reasonable notion of an ideal scheme is one that factorizes the parameters in a way that is equivalent to factoring the outputs of the layer; something better is impossible.
@@
$\|\overline{I}({\bm{\mathsfit{X}}})\overline{P}({\bm{\mathsfit{W}}})-\mathcal{T}^{\mathrm{SVD}}_{P}\bigl(\overline{I}({\bm{\mathsfit{X}}})\overline{P}({\bm{\mathsfit{W}}})\bigr)\|_{F}=\|\overline{I}({\bm{\mathsfit{X}}})\overline{P}({\bm{\mathsfit{W}}})-\overline{I}({\bm{\mathsfit{X}}})\mathcal{T}_{P}\bigl(\overline{P}({\bm{\mathsfit{W}}})\bigr)\|.$
Recall that each group is factorized separately. In particular, SVD-LLM (Wang et al., 2025c) introduces a particular way of “talking in terms of activation distortion” for linear layers, and it is optimal in the sense of Definition 2. A more general version of their low-rank activation-aware projection from the lens of our framework can be defined as follows. For any $(\overline{O},\overline{I},\overline{P})$-expressible function $f$, let
@@
For the layers and their corresponding $(\overline{O},\overline{I},\overline{P})$-tuples mentioned earlier, this corresponds, as discussed in previous examples, to two sequential sub-layers (e.g., two linear or convolutional layers) that yield computational gains with sufficiently low $P$. More details are given in Appendix B. Additionally, our scheme remains optimal, as claimed in Theorem 1.
@@
Recall that $B$ is the batch size (number of calibration samples) and $N$ the outer dimension (see the discussion below Definition 1). The main result of Wang et al. (2025c) follows as a corollary.
@@
4.2 Computing Whitening Matrices
We compute the whitening matrices with an eigendecomposition of the transformed features. Suppose we are given a group $g$. Let ${\bm{X}}=\overline{I}({\bm{\mathsfit{X}}})_{g}\in\mathbb{R}^{N\times I}$. We can then compute the eigenvectors and eigenvalues of $\frac{1}{N}{\bm{X}}^{T}{\bm{X}}$: $\bm{V}$} and $\bm{\Lambda}$}, respectively, and set
@@ 4.2 Computing Whitening Matrices
This procedure is applied (in vectorized form) independently for each group. Second moments can be accumulated incrementally over minibatches, so the method is memory-efficient. Additional implementation details are provided in Appendix C.
@@ 4.3 Budgeted Rank Allocation
Apart from offering output distortion guarantees, this framework is interpretable in the sense that we can measure the output distortion without additional steps (see Theorem 2). Can that be used to efficiently choose the rank to truncate each layer to?
@@
We now turn to the question of how to select the ranks to retain in each layer, $\{P_{l}\}_{l=1}^{L}$. The naive option is to choose a uniform percentage of rank per layer (Wang et al., 2025c). A more reasonable option, used in the past for the traditional SVD scheme, is the energy-based singular value pruning criterion (Yang et al., 2020; Liebenwein et al., 2021). While more expensive approaches have been explored, such as training-based rank learning (Idelbayev & Carreira-Perpiñán, 2020) or NAS-based search (Yu & Bouganis, 2022), our main goal is to find a scalable and efficient method. We can define the analog of the energy metric in our setting as follows (it reads as the energy retained on layer $l$ when keeping up to rank $P$):
$E_{l}(P)=1-\frac{\|f_{l}({\bm{\mathsfit{X}}}^{l-1};{\bm{\mathsfit{W}}}^{l})-f_{l}({\bm{\mathsfit{X}}}^{l-1};\mathcal{T}^{\mathrm{AA}}_{P}\bigl({\bm{\mathsfit{W}}}^{l}\bigr))\|^{2}_{F}}{\|f_{l}({\bm{\mathsfit{X}}}^{l-1};{\bm{\mathsfit{W}}}^{l})\|^{2}_{F}}=\frac{\sum_{g=1}^{G}\sum_{i=1}^{P}\bm{\sigma}_{g,i}^{2}}{\sum_{g=1}^{G}\sum_{i=1}^{U}\bm{\sigma}_{g,i}^{2}},$
where $l$ denotes the layer index, $U$ the total number of singular values (per group) and $P$ the number retained (per group). Energy-based selection then consists of picking the lowest rank such that some energy threshold (manually defined by the user) is retained. We propose instead to solve an optimization problem that maximizes the total retained energy globally:
@@
We solve this limitation by using a Lagrangian relaxation of the problem, which is linear in time w.r.t. $\sum_{l}U_{l}$, with a constant multiplicative cost associated with a hyperparameter $I$ (which stands for “number of iterations”). In practice, using 300 iterations yields excellent results while taking less than 0.2 seconds, even for the larger models we experimented with. We provide additional details in Section D.1.

### v2

@@ 1 Introduction
Here, we propose a unified framework that builds on Wang et al. (2025c) and is compatible with a broad class of layers. Moreover, we introduce a scalable rank allocator that determines how much to compress each layer based on a Lagrangian relaxation. Together, these form BALF, a complete and general fine-tuning-free compression pipeline. Our main contributions are:
@@
  • Extensive evaluation. Across vision models from ResNet-20 (He et al., 2016) on CIFAR-10 (Krizhevsky, 2009) to ResNeXt-101 (Xie et al., 2017) and ViT-B/16 (Dosovitskiy et al., 2021) on ImageNet (Deng et al., 2009), our method achieves strong accuracy-compression trade-offs in the fine-tuning-free regime.
@@ 2 Related Work
Deep learning model compression has been extensively studied. The principal categories of compression techniques include quantization, which employs low-bit storage and operations (Jacob et al., 2018; Frantar et al., 2023; Gholami et al., 2021); pruning, which involves the elimination of parameters or parameter groups from the model (Hoefler et al., 2021; Frantar & Alistarh, 2022); and factorization, which divides model layers into smaller sub-layers to lower overall memory usage and computational demands (Jaderberg et al., 2014; Ben Noach & Goldberg, 2020). This paper focuses on the factorization approach.
@@
Existing factorization approaches often require fine-tuning and/or compute-intensive search (Idelbayev & Carreira-Perpiñán, 2020; Yu & Bouganis, 2023). The impossibility of running these on common hardware motivates the need for lighter methods. For example, pruning methods that explicitly aim to be efficient in this regard include Narshana et al. (2023); Murti et al. (2023); Chen et al. (2024); Zhang et al. (2024); Wang et al. (2025a), whereas fine-tuning-free factorization methods are less common—for instance, Yu & Bouganis (2023). Liebenwein et al. (2021); Idelbayev & Carreira-Perpiñán (2020) also report some results in the fine-tuning-free regime.
@@ 3 Preliminaries and Notation
Given a matrix ${\bm{A}}\in\mathbb{R}^{M\times N}$, we denote its SVD by ${\bm{A}}={\bm{U}}{\bm{\Sigma}}{\bm{V}}^{T}$, where ${\bm{U}}\in\mathbb{R}^{M\times M}$ and ${\bm{V}}\in\mathbb{R}^{N\times N}$ are orthogonal matrices, and ${\bm{\Sigma}}=\operatorname{diag}(\bm{\sigma};(M,N))$ is a diagonal matrix with its singular values. Another concept of interest is the $P$-truncated SVD, which is defined by taking only the $P$ largest singular values and their associated singular vectors. In particular, we define $\mathcal{T}^{\mathrm{SVD}}_{P}\bigl({\bm{A}}\bigr)={\bm{U}}_{:,:P}{\bm{\Sigma}}_{:P,:P}({\bm{V}}_{:,:P})^{T}$. It can be thought of as an operator that projects a matrix to its rank-$P$ approximation via SVD.
@@
Data whitening (Kessy et al., 2018) (also termed data sphering) is a widely used technique in machine learning and statistics. Here, we are interested in a special case called uncentered whitening. This process transforms data so that its second-moment matrix becomes the identity matrix. Formally, we define an uncentered whitening matrix (UWM) ${\bm{M}}\in\mathbb{R}^{I\times I}$ for the data ${\bm{X}}\in\mathbb{R}^{N\times I}$ as a matrix such that ${\bm{M}}^{T}{\bm{X}}^{T}{\bm{X}}{\bm{M}}=N\overline{{\bm{I}}}_{R}$, where $\overline{{\bm{I}}}_{R}\in\mathbb{R}^{I\times I}$ is the identity on the first $R$ coordinates and zero elsewhere, and $R$ is the rank of $\bm{X}$}. We further restrict the definition by requiring that $\operatorname{row}({\bm{X}})=\operatorname{range}({\bm{M}})$, which in turn forces the last IRI-R columns of $\bm{M}$} to be zero and makes ${\bm{X}}={\bm{X}}{\bm{M}}{\bm{M}}^{+}$; this assumption and these consequences will be used in later results.
@@
We refer to general tensors as $\bm{\mathsfit{X}}$}. When some tensor is known to be a matrix or a vector, we may use $\bm{X}$} and $\bm{x}$}, respectively. For matrices, we denote the Moore-Penrose inverse (also termed the pseudoinverse) of $\bm{M}$} by $\bm{M}$$^{+}$, and the square root of the pseudoinverse by ${\bm{M}}^{+1/2}$.
Deep learning frameworks, such as PyTorch (Ansel et al., 2024), usually include “reshape” and “permute” operators, denoted in this work by $\mathsf{reshape}\left({\bm{\mathsfit{X}}};(s_{1},\dots,s_{n})\right)$ and $\mathsf{permute}\left({\bm{\mathsfit{X}}};(\pi_{1},\dots,\pi_{n})\right)$, respectively.
@@
We denote the identity function by $\mathrm{id}$. Moreover, when discussing convolutional layers, $B$ will denote the batch size; $H_{x},W_{x},C_{x}$ the height, width, and channels of the input (when $x=i$) and output (when $x=o$); and $H_{k},W_{k}$ the spatial size of the kernel. In this context, we also refer to $G$ as the number of groups in the convolutional layer (Xie et al., 2017).
@@ 4 Our Framework
$G$ stands for the number of groups. In the special case where $G=1$, we may simply refer to matrices instead of batches containing a single matrix. Note that in the preceding definition, matrix multiplication is batched, i.e., $G$ independent matrix multiplications are carried out. The necessity of batch semantics in the group dimension will become apparent later. This condition is satisfied by popular layers. In what follows, layers are assumed to operate in batched mode, with $B$ denoting the batch size. We refer to $N$ as the outer dimension.
@@
In some architectures (e.g., transformers), linear layers receive inputs with multiple leading independent dimensions—for example, a tensor of shape $(B,L,D)$. In those cases, we flatten the leading dimensions, reshaping the input to $(BL,D)$ and restoring the original shape afterward. We now turn our attention to convolutional layers. Although an ungrouped convolution is a special case of a (possibly grouped) convolution, we present it separately for clarity.
@@
where $\mathsf{im2col}:\mathbb{R}^{B\times C_{i}\times H_{i}\times W_{i}}\to\mathbb{R}^{B\times H_{o}W_{o}\times C_{i}H_{k}W_{k}}$ maps a batch of 2D signals with $C_{i}$ channels into a matrix (see Chellapilla et al. (2006)). The PyTorch equivalent is called $\mathsf{unfold}$ [^1] https://docs.pytorch.org/docs/stable/generated/torch.nn.Unfold.html. It encapsulates the convolution hyperparameters—kernel size, stride, padding, and dilation—but not the number of groups. For brevity, we omit the encapsulated hyperparameters from the notation.
On the other hand, a grouped convolution can be viewed as $G$ separate, parallel convolutions. Each one takes a disjoint slice of $C_{i}/G$ input channels and produces $C_{o}/G$ output channels. Concatenating the $G$ outputs recovers the full $C_{o}$ channels (Krizhevsky et al., 2012; Xie et al., 2017). This also fits our framework.
@@
Additionally, Example 2 and Example 3 have trivial extensions to higher-dimensional convolutions, achieved by replacing $\mathsf{im2col}$ with its higher-dimensional counterparts.
@@
[…]
@@
$\mathcal{T}^{D}_{P}\bigl(\overline{P}({\bm{\mathsfit{W}}})_{g}\bigr)=\operatorname{arg}\,\underset{\operatorname{rank}(\overline{P}(\widehat{{\bm{\mathsfit{W}}}})_{g})\leq P}{\min}\;D(\overline{P}({\bm{\mathsfit{W}}})_{g},\overline{P}(\widehat{{\bm{\mathsfit{W}}}})_{g}),$ (2)
where $D$ is some penalty function, and $\widehat{{\bm{\mathsfit{W}}}}$ is the approximate tensor. Traditional SVD-based decomposition corresponds to the case where $D$ is the Frobenius norm of the (reshaped) parameters, i.e., it penalizes distance in parameter space (see Equation 1). We use $\mathcal{T}^{D}_{P}\bigl({\bm{\mathsfit{W}}}\bigr)$ to also denote batched computation, where each group’s matrix is decomposed individually. As before, we omit $\overline{P}$ to simplify notation.
@@ 4.1 Activation-Aware Low-Rank Projections
In order to answer that question, we first need to define what “best” should mean. Without loss of generality, fix a group $g$. Clearly, $\operatorname{rank}\left(\overline{I}({\bm{\mathsfit{X}}})\mathcal{T}_{P}\bigl({\bm{\mathsfit{W}}}\bigr)\right)\leq P$. Moreover, by the EYM theorem, the best rank-$P$ approximation of the output of $f$ in terms of the Frobenius norm is $\mathcal{T}^{\mathrm{SVD}}_{P}\bigl(\overline{I}({\bm{\mathsfit{X}}})\overline{P}({\bm{\mathsfit{W}}})\bigr)$. Hence, a reasonable notion of an ideal scheme is one that factorizes the parameters in a way that is equivalent to factorizing the outputs of the layer; something better is impossible.
@@
$\|\overline{I}({\bm{\mathsfit{X}}})\overline{P}({\bm{\mathsfit{W}}})-\mathcal{T}^{\mathrm{SVD}}_{P}\bigl(\overline{I}({\bm{\mathsfit{X}}})\overline{P}({\bm{\mathsfit{W}}})\bigr)\|_{F}=\|\overline{I}({\bm{\mathsfit{X}}})\overline{P}({\bm{\mathsfit{W}}})-\overline{I}({\bm{\mathsfit{X}}})\mathcal{T}_{P}\bigl(\overline{P}({\bm{\mathsfit{W}}})\bigr)\|_{F}.$
Recall that each group is factorized separately. Note that a scheme is optimal if and only if it satisfies Equation 2 with the distance function set to the output distortion. In particular, SVD-LLM (Wang et al., 2025c) introduces a specific way of “talking in terms of activation distortion” for linear layers, and it is optimal in the sense of Definition 2. A more general version of their low-rank activation-aware projection method, from the lens of our framework, can be defined as follows. For any $(\overline{O},\overline{I},\overline{P})$-expressible function $f$, let
@@
For the layers and their corresponding $(\overline{O},\overline{I},\overline{P})$-tuples mentioned earlier, this corresponds, as discussed in previous examples, to two sequential sub-layers (e.g., two linear or convolutional layers) that yield computational gains with sufficiently low $P$. More details are given in Appendix A. Additionally, our scheme remains optimal, as claimed in Theorem 1.
@@
Recall that $B$ is the batch size (number of calibration samples), and $N$ is the outer dimension (see the discussion below Definition 1). The main result of Wang et al. (2025c) follows as a corollary.
@@
4.2 Computing the Whitening Matrices
We compute the whitening matrices using an eigendecomposition of the transformed features. Suppose we are given a group $g$. Let ${\bm{X}}=\overline{I}({\bm{\mathsfit{X}}})_{g}\in\mathbb{R}^{N\times I}$. We can then compute the eigenvectors and eigenvalues of $\frac{1}{N}{\bm{X}}^{T}{\bm{X}}$: $\bm{V}$} and $\bm{\Lambda}$}, respectively, and set
@@ 4.2 Computing the Whitening Matrices
This procedure is applied (in vectorized form) independently for each group. Second moments can be accumulated incrementally over minibatches, so the method is memory-efficient. Additional implementation details are provided in Appendix B.
@@ 4.3 Budgeted Rank Allocation
Apart from offering output distortion guarantees, this framework is interpretable in the sense that we can measure the output distortion without additional steps (see Theorem 2). Can that be used to efficiently choose the rank to which each layer should be truncated?
@@
We now turn to the question of how to select the ranks to retain in each layer, $\{P_{l}\}_{l=1}^{L}$. The naive option is to choose a uniform percentage of rank per layer (Wang et al., 2025c). A more reasonable option, used in the past for the traditional SVD scheme, is the energy-based singular value pruning criterion (Yang et al., 2020; Liebenwein et al., 2021). While more expensive approaches have been explored, such as training-based rank learning (Idelbayev & Carreira-Perpiñán, 2020; Qinsi et al., 2025) or NAS-based search (Yu & Bouganis, 2023), our main goal is to find a scalable and efficient method. We can define the analog of the energy metric in our setting as follows (it reads as the energy retained in layer $l$ when keeping up to rank $P_{l}$):
$E_{l}(P_{l})=1-\frac{\|f_{l}({\bm{\mathsfit{X}}}^{l-1};{\bm{\mathsfit{W}}}^{l})-f_{l}({\bm{\mathsfit{X}}}^{l-1};\mathcal{T}^{\mathrm{AA}}_{P_{l}}\bigl({\bm{\mathsfit{W}}}^{l}\bigr))\|^{2}_{F}}{\|f_{l}({\bm{\mathsfit{X}}}^{l-1};{\bm{\mathsfit{W}}}^{l})\|^{2}_{F}}=\frac{\sum_{g=1}^{G}\sum_{i=1}^{P_{l}}\bm{\sigma}_{g,i}^{2}}{\sum_{g=1}^{G}\sum_{i=1}^{U_{l}}\bm{\sigma}_{g,i}^{2}},$
where $l$ denotes the layer index, $U_{l}$ the total number of singular values (per group), and $P_{l}$ the number retained (per group). Energy-based selection then consists of picking the lowest rank such that a certain energy threshold (manually defined by the user) is retained. We propose instead to solve an optimization problem that maximizes the total retained energy globally:
@@
We solve this limitation by using a Lagrangian relaxation of the problem, whose time complexity is $O((I+D+1)\sum_{l}U_{l})$, where $I$ stands for “number of iterations” and $D$ depends on the network. In practice, using 300 iterations yields excellent results while taking less than 0.2 seconds, even for the larger models with which we experimented. We provide additional details in Section C.1.